# A Comprehensive Reanalysis of K2-18 b’s JWST NIRISS+NIRSpec Transmission Spectrum

### v1

@@ 6.3 On Establishing Robust and Reliable
Our revised atmospheric composition for K2-18 b demonstrates the critical importance of using multiple data reductions and multiple retrieval codes to establish which molecular detections are robust. In particular, our non-detection of CO2 stands in stark contrast to the finding of a robust detection reported by Madhusudhan et al. (2023). However, this is not the first time that reported molecular detections in K2-18 b’s atmosphere have proven unreliable. Pre-JWST, multiple studies reported detections of H2O vapor from HST WFC3 transmission spectroscopy at compelling statistical significances: 3.6 (Tsiaras et al., 2019), 3.9 (Benneke et al., 2019), and 3.3 (Madhusudhan et al., 2020). We now know from the JWST spectrum of K2-18 b that the spectral feature seen by WFC3 was, in fact, CH4 masquerading as H2O (likely due to WFC3 edge effects near 1.7 ). One naturally wonders how so many retrieval ‘detections’ could prove unreliable. Thus, we examine here how future retrieval analyses can safeguard against overstating the statistical evidence of a molecular detection.
@@
First, our grid of retrievals demonstrates that detection significances naturally possess their own uncertainty due to data-level assumptions. Figure 5 clearly illustrates a spread of CH4 significances (clustered around ) as a function of the data reduction code and the degree of spectral binning pre-retrieval. We propose that visualizations such as Figure 5 are a valuable tool to intuitively establish if a molecular detection is robust (as with CH4) or whether there is no reliable statistical evidence (as with CO2 and DMS). An additional value of Figure 5 is the clear representation of lost information when retrieving NIRISS spectra at rather than , resulting in a significantly lower CH4 detection significance, which suggests is a safe minimum resolution for inferring atmospheric properties from JWST transmission spectra. Our first takeaway is thus that the term ‘robust’ must be reserved for molecular detections that have been explicitly demonstrated to hold across multiple data reduction and retrieval codes.
@@
Second, we argue that the retrieval literature tends to present overconfident detection significances and inconsistent terminology from Bayesian model comparisons. When establishing if a molecule is present in an atmosphere, the common practice is to conduct a Bayesian model comparison between a retrieval model containing the molecule in question and one with it removed, calculate the Bayes factor, and then convert the Bayes factor into an ‘equivalent detection significance’ (using the inverse of Equation 27 in Trotta 2008). This approach was first introduced in the exoplanet retrieval context by Benneke & Seager (2013) and has been widely used in retrieval codes since (e.g., Waldmann et al., 2015; MacDonald & Madhusudhan, 2017; Welbanks & Madhusudhan, 2021). This procedure often provides useful insights, provided that: (i) both the Bayes factor and equivalent significance are provided, and (ii) well-defined terminology is adopted (e.g. the Jeffreys’ scale; see Trotta 2008). Not providing Bayes factors can produce overconfidence in a result due to common intuition for the meaning of an ‘’ result. For example, a Bayes factor of 3 corresponds to an equivalent significance of , such that readers with frequentist training may construe the detection confidence to be a 95% result, when the model containing the molecule is actually only favored at a 3:1 odds ratio. Similarly, a claim of a ‘detection’ from a Bayesian model comparison is only a Bayes factor of 21. Trotta (2008) highlights this issue by recommending that an ‘’ result should be thought of as an ‘’ result to calibrate the frequentist measure for what the Bayes factor represents. Using the same logic, the ‘ detection’ of CO2 in K2-18 b’s atmosphere reported by Madhusudhan et al. (2023) (without quoted Bayes factors) should be better thought of as a result, for which the term ‘detection’ is too strong. It is for this reason that we report all Bayes factors 3 (equivalent to ) as non-detections in Figure 5, since in reality these represent odds ratios closer to a ‘’ result via the ‘’ calibration. Clearly, the conflation of a equivalent significance from Bayesian model comparisons as being equivalent to a 99.7% detection confidence has proven especially problematic in the case of K2-18 b — as illustrated by the reported ‘detections’ of both H2O (Benneke et al., 2019; Tsiaras et al., 2019; Madhusudhan et al., 2020) and CO2 Madhusudhan et al. (2023) proving unreliable for K2-18 b following new observations and/or a revised retrieval analysis.
@@
| Molecule | Our Classification | Their Classification | Our Classification | ||||||
| CH4 | 86– | 3.4–5.6 | 2.4–4.6 | robust detection | 11,000–44,000† | 4.7–5.0 | 3.7–4.0 | robust detection | detection‡ |
| CO2 | 0.6–4.2 | no evidence | 18–40† | 2.9–3.2 | 1.9–2.2 | robust detection | moderate evidence | ||
| DMS | 0.6–3.9 | no evidence | 1–5† | potential signs | no evidence | ||||
@@
Note. — is the Bayes factor between models with and without molecule ‘i’. Our classification uses the following terms: ‘no evidence’ (), ‘weak evidence’ (), ‘moderate evidence’ (), ‘strong evidence’ (), and ‘detection’ (). We reserve ‘robust detection’ for cases where a grid of retrievals across multiple data-level and model-level configurations all find . For cases with some retrievals in a grid reside in one category but others fall in another category, the category label corresponding to most of the Bayesian model comparisons takes precedence over outliers. †Bayes factors were not provided by Madhusudhan et al. (2023), so we compute them from their quoted detection significances. The range of Bayes factors and significances quoted for Madhusudhan et al. (2023) correspond to their no offset, one offset, and two offsets cases. ‡The CH4 detection from Madhusudhan et al. (2023) does not qualify for the ‘robust’ label under our classification (even though ), because their three retrievals only considered a single data reduction and retrieval code. We exclude NIRISS data from our quoted CH4 Bayes factor and significance ranges, since data with incurs degeneracies that hinder CH4 detectability (see Section 4.2.1 and Figure 5). Since the formula mapping Bayes factors to significances breaks down for low Bayes factors (), we quote only the highest significance for CO2 and DMS.
We propose a straightforward classification scheme for communicating atmospheric retrieval results to mitigate these issues. Our starting point is the widely used Jeffreys’ scale, which speaks only to Bayes factors and not equivalent significances. We propose the following terms for low-statistical significances: ‘no evidence’ (Bayes factor 3), ‘weak evidence’ (3 Bayes factor 12), ‘moderate evidence’ (12 Bayes factor 150), and ‘strong evidence’ (150 Bayes factor 600). We propose that a Bayes factor of 600 should represent the terminology transition from ‘evidence’ to ‘detection’, since this Bayes factor corresponds to an odds ratio of 99.8% — close to the traditional understanding of what a 3 result represents. Where ‘equivalent detection significances’ () are reported, we advocate for also reporting the significance. Finally, we propose that the term ‘robust detection’ be reserved for when the vast majority of Bayesian model comparisons for a molecule result in detections (i.e. Bayes factor 600 across a range of data-level and model-level configurations). Consequently, under our proposed classification scheme, a detection can only be considered robust if it is confirmed with multiple data reduction and retrieval codes.
@@
Table 5 demonstrates the application of our molecular detection classification scheme to K2-18 b transmission spectra retrievals. We compare the Bayes factors, equivalent significances (‘’), and calibrated significances (‘’) from Madhusudhan et al. (2023) and our study. For the statistics from Madhusudhan et al. (2023), we also compare the detection terminology used in that study with our proposed classification under the scheme outlined above. First, let us consider the evidence for CH4 in K2-18 b’s atmosphere. Figure 5 shows that 33/40 Bayesian model comparisons with have , corresponding to Bayes factors 600, which we regard as sufficient to ascribe CH4 as a ‘robust detection’ (only the FIREFLy NIRISS data results in slightly lower significances). The CH4 significances from Madhusudhan et al. (2023) meet our threshold for ‘detection’, but do not qualify as a ‘robust detection’ under our classification due to the use of a single data reduction code and a single retrieval code. Second, we see that the statistics reported for CO2 by Madhusudhan et al. (2023) would be downgraded from ‘robust detection’ to ‘moderate evidence’ under our scheme, while our retrievals are classified as ‘no evidence’ for CO2 (see Figure 5). The CO2 classification demonstrates the utility of our proposed scheme, since the terminology readily discriminates between the reliable CH4 inference and the significantly lower evidence for CO2 in the JWST data from Madhusudhan et al. (2023) (see also Appendix A). Finally, we see that our classification scheme reports ‘no evidence’ for DMS for both our retrieval results and those from Madhusudhan et al. (2023). We highlight that the ‘’ column in Table 5 intuitively maps to the corresponding evidence classification for each molecule. We therefore advocate for adopting this more measured terminology for Bayesian model comparisons to avoid overstating the statistical evidence for molecular detections from JWST spectra of exoplanet atmospheres.
6.4 Upcoming JWST Observations of K2-18 b
@@
JWST will continue to observe spectra of K2-18 b, offering deeper insights into its atmospheric composition and internal structure. Two approved JWST programs are undertaking additional reconnaissance of K2-18 b’s atmosphere: Program GO-2372 (PI: Renyu Hu) and Program GO-2722 (PI: Nikku Madhusudhan). We briefly discuss how our results relate to potential outcomes from these upcoming observations.
@@
Program GO-2372 is in the process of observing 4 additional transits with NIRSpec G395H (with 2 pending at the time of writing) and 2 transits with NIRSpec G235H. These observations will significantly improve the observational precision over the 1.7–5.2 wavelength range, offering additional sensitivity to the allowed abundances of molecules including CH4, CO2, DMS, CO, H2O, NH3, and HCN. The improved precision of these observations may allow the detection of CO2, which would be expected for a mini-Neptune with a slightly lower abundance than our present CO2 upper limit (see Figures 7 and 9). Therefore, a detection of CO2 in new multi-transit JWST data is not at odds with the non-detection of CO2 from the single NIRISS SOSS and NIRSpec G395H transits we report here.
Program GO-2722 has observed two transits with MIRI LRS, which will offer the first transmission spectrum of K2-18 b from 5–12. These MIRI observations are anticipated to be especially sensitive to an array of sulfur-bearing species, including DMS and CS2 (Tsai et al., 2024). However, the low signal-to-noise ratio in MIRI LRS transmission spectra at wavelengths near 10 (e.g., Powell et al., 2024), where DMS absorption has a strong absorption feature, may preclude any significant detections of DMS from MIRI LRS spectra alone. Therefore, any claim of a DMS detection from MIRI LRS observations must rely on additional information content from the shorter near-infrared wavelengths (i.e. NIRSpec G395H). Since our results demonstrate that the present JWST observations of K2-18 b do not provide any reliable statistical evidence for DMS, it is unlikely that these MIRI observations can further elucidate whether or not DMS is present in K2-18 b’s atmosphere. We also stress that any inference of DMS must be considered in light of the poor availability of opacity data for this molecule, as the DMS cross section from Sharpe et al. (2004), used in all K2-18 b retrievals, has no temperature and pressure dependence. A variety of other chemical species with C-S bonds should also be considered when modeling MIRI LRS transmission spectra to cover degenerate interpretations. Consequently, we regard further observations of K2-18 b, beyond the two Programs described above, as being necessary to achieve the necessary sensitivity to unambiguously detect sulfur-bearing molecules in K2-18 b’s atmosphere.
@@ Acknowledgments
We thank Adina Feinstein, Duncan Christie, Emma Esparza-Borges, Evert Nasedkin, and Jake Taylor for helpful comments. We thank Hamish Innes for an illuminating discussion on the supercritical ocean scenario. This work is based in part on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The specific observations analyzed can be accessed via https://doi.org/10.17909/3ds1-8z15 (catalog DOI:10.17909/3ds1-8z15). S.P.S. is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2139757. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation. R.J.M. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51513.001, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS 5-26555. S.-M.T. acknowledges support from NASA Exobiology grant No. 80NSSC20K1437 and the University of California, Riverside. M.R. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC). T.J.B. acknowledges funding support from the NASA Next Generation Space Telescope Flight Investigations program (now JWST) via WBS 411672.07.04.01.02. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has made use of NASA’s Astrophysics Data System.

### v2

@@ 6.2 On Establishing Robust and Reliable
Our revised atmospheric composition for K2-18 b demonstrates the critical importance of using multiple data reductions and multiple retrieval codes to establish which molecular detections are robust. In particular, our non-detection of CO2 stands in stark contrast to the finding of a robust detection reported by Madhusudhan et al. (2023). However, this is not the first time that reported molecular detections in K2-18 b’s atmosphere have proven unreliable. Pre-JWST, multiple studies reported detections of H2O vapor from HST WFC3 transmission spectroscopy at compelling statistical significances: 3.6 (Tsiaras et al., 2019), 3.9 (Benneke et al., 2019a), and 3.3 (Madhusudhan et al., 2020). We now know from the JWST spectrum of K2-18 b that the spectral feature seen by WFC3 was, in fact, CH4 masquerading as H2O (likely due to WFC3 edge effects near 1.7 ). Here we discuss here how future retrieval analyses can safeguard against overstating the statistical evidence of a molecular detection.
@@
First, our grid of retrievals demonstrates that detection significances naturally possess their own uncertainty due to data-level assumptions. Figure 5 clearly illustrates a spread of CH4 significances (clustered around ) as a function of the data reduction code and the degree of spectral binning pre-retrieval. We propose that visualizations such as Figure 5 are a valuable tool to intuitively establish if a molecular detection is robust (as with CH4) or whether there is no reliable statistical evidence (as with CO2 and DMS). An additional value of Figure 5 is the clear representation of lost information when retrieving NIRISS spectra at rather than , resulting in a significantly lower CH4 detection significance, which suggests is a safe minimum resolution for inferring atmospheric properties from JWST transmission spectra. Our first takeaway is thus that the term ‘robust’ must be reserved for molecular detections that hold across multiple data reduction and retrieval codes.
@@
Second, we note that the retrieval literature tends to present overconfident detection significances and inconsistent terminology from Bayesian model comparisons. When establishing if a molecule is present in an atmosphere, the common practice is to conduct a Bayesian model comparison between a retrieval model containing the molecule in question and one with it removed, calculate the Bayes factor, and then convert the Bayes factor into an ‘equivalent detection significance’ (using the inverse of Equation 27 in Trotta 2008). This approach was first introduced in the exoplanet retrieval context by Benneke & Seager (2013) and has been widely used in retrieval codes since (e.g., Waldmann et al., 2015; MacDonald & Madhusudhan, 2017; Welbanks & Madhusudhan, 2021). This procedure often provides useful insights, provided that: (i) both the Bayes factor and equivalent significance are provided, and (ii) well-defined terminology is adopted (e.g. the Jeffreys’ scale; see Trotta 2008). Not providing Bayes factors can produce overconfidence in a result. For example, a Bayes factor of 3 corresponds to an equivalent significance, according to the approach from Benneke & Seager (2013), of , when this model is actually only favored at a 3:1 odds ratio. Similarly, a claim of a ‘detection’ from a Bayesian model comparison is only a Bayes factor of 21. Recently, Kipping & Benneke (2025) highlighted this point and suggested that a more intuitive detection significance is given by equating the false positive detection probability with the value (i.e. ) and deriving . This results in values of that are approximately 1 lower than the relation from Benneke & Seager (2013) (see also Trotta 2008). Using this logic, the ‘ detection’ of CO2 in K2-18 b’s atmosphere reported by Madhusudhan et al. (2023) (without quoted Bayes factors) should be better thought of as a result, for which the term ‘detection’ is too strong. It is for this reason that we report all Bayes factors 3 (equivalent to ) as non-detections in Figure 5.
@@
| Molecule | Jeffreys’ Scale | Their Classification | Jeffreys’ Scale | ||||||
| CH4 | 12– | 2.7–5.6 | 1.8–4.9 | moderate– | 11,000–44,000† | 4.7–5.0 | 3.9–4.2 | robust detection | strong evidence |
| strong evidence | |||||||||
| CO2 | 0.6–4.2 | no evidence | 18–40† | 2.9–3.2 | 1.9–2.3 | robust detection | moderate evidence | ||
| DMS | 0.6–3.9 | no evidence | 1–5† | potential signs | no evidence– | ||||
| weak evidence | |||||||||
@@
Note. — is the Bayes factor between models with and without molecule ‘i’. uses the relation from Benneke & Seager (2013), while is derived from equating ) (Kipping & Benneke, 2025). The Jeffreys’ scale uses the following terms: ‘no evidence’ (), ‘weak evidence’ (), ‘moderate evidence’ (), and ‘strong evidence’ (). †Bayes factors were not provided by Madhusudhan et al. (2023), so we compute them from their quoted detection significances. The range of Bayes factors and significances quoted for Madhusudhan et al. (2023) correspond to their no offset, one offset, and two offsets cases. Our lowest Bayes factor for CH4 corresponds to NIRISS data, which hinders the CH4 detectability (see Section 4.2.1 and Figure 5), but for data the minimum CH4 Bayes factor is 86 (). Since the formula mapping Bayes factors to breaks down for low Bayes factors (), we quote only the highest significance for CO2 and DMS.
@@
In Table 5, we summarize and compare the statistical evidence for CH4, CO2, and DMS in K2-18 b’s atmosphere from our study with Madhusudhan et al. (2023). We present the Bayes factors, the ‘classical’ detection significance (following the scheme from Benneke & Seager 2013) and the ‘intuitive’ detection significance (Kipping & Benneke, 2025). First, we consider the evidence for CH4 in K2-18 b’s atmosphere. Figure 5 shows that 33/40 Bayesian model comparisons with have — or — which corresponds to ‘strong evidence’ on the Jeffreys’ scale (only the FIREFLy NIRISS data results in slightly lower significances). At this level of model preference, the term ‘detection’ is clearly warranted for CH4 in both our study and Madhusudhan et al. (2023). Second, we see that the statistics reported for CO2 by Madhusudhan et al. (2023) should be downgraded from ‘robust detection’ to ‘moderate evidence’ according to the Jeffreys’ scale, while our retrieval model comparisons are classified as ‘no evidence’ for CO2 (see Figure 5). These results highlight the reliable and reproducible CH4 inference, but show that the JWST data from Madhusudhan et al. (2023) do not provide sufficient evidence supporting a detection of CO2 (see also Appendix A). Finally, we see that our retrievals yield ‘no evidence’ for DMS, while the results from Madhusudhan et al. (2023) correspond to ‘no evidence’ or ‘weak evidence’ at most when using the Jeffreys’ scale. We therefore advocate for reporting more conservative statistical measures for detection significances from Bayesian model comparisons (e.g. , see Kipping & Benneke 2025), alongside exploring the sensitivity of detections to specific data regions (e.g. Welbanks et al., 2023), to avoid overstating the statistical evidence of molecules from JWST spectra of exoplanet atmospheres.
6.3 Compatibility Between Near-infrared and Mid-infrared Transmission Spectra of K2-18 b
Here, we briefly examine the implications of our non-detections of DMS and CO2 with the mid-infrared MIRI LRS spectrum of K2-18 b recently presented by Madhusudhan et al. (2025). In Figure 14 (upper panel), we show our best-fitting transmission spectrum to our FIREFLy low-resolution data (NIRISS SOSS at , NIRSpec G395H at ) extrapolated to the mid-infrared. We see that this extrapolated model — which contains only a minor CH4 feature and no DMS absorption in the mid-infrared — provides a reasonable fit to the MIRI LRS JExoRES data from Madhusudhan et al. (2025), consistent with there being no significant spectral features in the mid-infrared (Taylor, 2025; Welbanks et al., 2025). However, the same is not true when extrapolating the transmission spectrum corresponding to the median retrieved atmospheric properties reported by Madhusudhan et al. (2025) backward to the near-infrared. We show in Figure 14 (lower panel) that the DMDS features reported by Madhusudhan et al. (2025) would produce significantly stronger absorption near 2.3 and 3.3 than is seen in the near-infrared data (or in the corresponding median retrieved models from Madhusudhan et al. 2023). This mismatch in scale heights is primarily driven by the MIRI LRS data favoring a temperature of 350 K, while the near-infrared data favors cooler temperatures of 240 K. We note that Luque et al. (2025) similarly found that simultaneous retrievals of the near-infrared and mid-infrared JWST data offer no significant evidence for DMS or DMDS in K2-18 b’s atmosphere.
6.4 Additional JWST Observations of K2-18 b
@@
JWST will continue to observe spectra of K2-18 b, offering deeper insights into its atmospheric composition and internal structure. In particular, Program GO-2372 (PI: Renyu Hu) is obtaining new near-infared observations of K2-18 b’s transmission spectrum. We briefly discuss how our results relate to potential outcomes from these upcoming observations.
@@
Program GO-2372 is in the process of observing 4 additional transits with NIRSpec G395H (with 2 pending at the time of writing) and 2 transits with NIRSpec G235H. These observations will significantly improve the observational precision over the 1.7–5.2 wavelength range, offering additional sensitivity to the allowed abundances of molecules including CH4, CO2, DMS, CO, H2O, NH3, and HCN. The improved precision of these observations may allow the detection of CO2, which would be expected for a mini-Neptune with a slightly lower abundance than our present CO2 upper limit (see Figure 9). Therefore, a detection of CO2 in new multi-transit JWST data would not be not at odds with the non-detection of CO2 from the single NIRISS SOSS and NIRSpec G395H transits we report here.
6.5 K2-18 b is Unlikely to Harbor a Water Ocean
@@ Acknowledgments
We thank the anonymous referee for their helpful comments. We also thank Adina Feinstein, Duncan Christie, Emma Esparza-Borges, Evert Nasedkin, and Jake Taylor for helpful comments. We thank Hamish Innes for an illuminating discussion on the supercritical ocean scenario. This work is based in part on observations made with the NASA/ESA/CSA JWST. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. The specific observations analyzed can be accessed via https://doi.org/10.17909/3ds1-8z15 (catalog DOI:10.17909/3ds1-8z15) on MAST (Madhusudhan, 2023). S.P.S. is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2139757. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author and do not necessarily reflect the views of the National Science Foundation. R.J.M. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51513.001, awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS 5-26555. S.-M.T. acknowledges support from NASA Exobiology grant No. 80NSSC20K1437 and the University of California, Riverside. M.R. acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC). T.J.B. acknowledges funding support from the NASA Next Generation Space Telescope Flight Investigations program (now JWST) via WBS 411672.07.04.01.02. This research was supported in part through computational resources and services provided by Advanced Research Computing at the University of Michigan, Ann Arbor. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This research has made use of NASA’s Astrophysics Data System.